### v2

@@ Abstract
Speculative decoding in large language models (LLMs) accelerates token generation by speculatively predicting multiple tokens cheaply and verifying them in parallel, and has been widely deployed. In this paper, we provide the first study demonstrating the privacy risks of speculative decoding. We observe that input-dependent patterns of correct and incorrect predictions can be leaked out to an adversary monitoring token generation times and packet sizes, leading to privacy breaches. By observing the pattern of correctly and incorrectly speculated tokens, we show that a malicious adversary can fingerprint queries and learn private user inputs with more than 90% accuracy across three different speculative decoding techniques – REST (almost 100% accuracy), LADE (up to 92% accuracy), and BiLD (up to 95% accuracy). We show that an adversary can also leak out confidential intellectual property used to design these techniques, such as data from data-stores used for prediction (in REST) at a rate of more than 25 tokens per second, or even hyper-parameters used for prediction (in LADE). We also discuss mitigation strategies, such as aggregating tokens across multiple iterations and padding packets with additional bytes, to avoid such privacy or confidentiality breaches.
@@ 1 Introduction
Large Language Models (LLMs) have transformed natural language processing (NLP), enabling machines to generate and understand human language at an unprecedented scale Vaswani et al. (2017); Devlin et al. (2019); Brown et al. (2020); Zhang et al. (2022); Le Scao et al. (2022); Touvron et al. (2023). LLMs typically generate text using auto-regressive decoding Touvron et al. (2023); OpenAI (2024a), where the generation happens serially and each token depends on all the previous ones. Unfortunately, this serial process causes a significant bottleneck in the LLM response latency Miao et al. (2024); Fu et al. (2024) and under-utilizes the available hardware-level parallelism, limiting token generation throughput and latency.
Speculative decoding Leviathan et al. (2023); Chen et al. (2023); Miao et al. (2024); Spector & Re (2023) addresses this problem without impacting model accuracy. It uses smaller models or heuristics such as retrieval or self-drafting to inexpensively generate tokens speculatively, which the larger target model verifies in parallel in a single iteration. By tuning the heuristics to maintain a high rate of correct speculations, such techniques provide 2x to 5x speedups in inference latency and throughput Xia et al. (2024).
However, the adoption of speculative techniques is not without risks. Speculative execution in CPUs  Burton (1985); Hennessy & Patterson (2012), which inspired speculative decoding, has led to considerable security vulnerabilities in processors, such as Spectre Kocher et al. (2019), Meltdown Lipp et al. (2018), and their numerous variants. These side-channel vulnerabilities exploit the side-effects of processor mis-speculations, such as timing variations introduced by mis-speculations, to leak private and confidential data that become accessible to a processor during mis-speculations. This raises the question, does speculative decoding in LLMs also introduce new risks to privacy?
In this paper, we provide the first study of the privacy and confidentiality risks of speculative decoding in LLMs, including leakage of private user inputs or confidential data.
@@
Q. How to fine-tune a large language model?   A.  F ine - t uning _a _large _language _model _involves _adjust ing _the _parameters _of _the _model _to _improve _its _performance _on _a _specific _task _or _down stream _task . _Here _are _some _steps _you _can _follow _to _fine - t une _a _large _language _model : <0x0A> <0x0A> 1 . _Cho ose _the _down stream _task : _Once _you _have _a _large _language _model , ...
(a) Speculative Decoding Example
(b) Tokens Per Autoregressive Iteration
@@
Consider a response generated by TinyLlama 1.1B Chat V1.0 Zhang et al. (2024a) with the state-of-the-art speculative decoding technique, Lookahead Decoding (LADE) Fu et al. (2024) in Figure 1(a), as an example. The red tokens are mis-speculated, requiring one auto-regressive iteration per token. Whereas, the blue tokens are correctly speculated, with up to $n$ tokens verified in a single iteration, as shown in Figure 1(b). Thus, in LLM applications that stream responses back to the user iteration by iteration, an adversary over the network, similar to recent attacks on LLMs Weiss et al. (2024), can identify network packets corresponding to each iteration and observe the packet size to deduce the number of tokens produced by each iteration and leak the pattern of correct and incorrect speculations dependent on the input.
Exploits. By observing speculation patterns, we demonstrate three possible exploits on speculative-decoding implementations, as shown in Figure 1(c).
First, we demonstrate a query fingerprinting attack by a network-based adversary. By using the speculation-pattern as a fingerprint for an prompt/response pair, the adversary can profile the fingerprints for different prompt/output pairs offline, and then for an unknown user query, measure the speculation-pattern, and compare it with the profiled fingerprints to leak out the user’s private prompt and response. On three different speculative decoding implementations (BiLD, REST, LADE) Fu et al. (2024); He et al. (2024); Kim et al. (2023), we show that one can identify user inputs using fingerprints with accuracy of up to $\sim$100for REST, 92% for LADE, and 95% for BiLD with just 30 training samples. When the prompts to be leaked are just approximately similar to profiled prompts, the accuracy is still within 20% to 40% accuracy (much better than random guessing probability of 2%).
Second, we show that an adversary capable of crafting malicious inputs can leak confidential data or hyper-parameters that may be used for generating predictions in speculative decoding techniques. If predictions are generated via retrieval from a data-store (REST) He et al. (2024), we show that an adversary can craft inputs and monitor which output tokens are correctly predicted, and leak out the data-store at the rate of >25 tokens per second. When speculative decoding generates predictions using self-drafting (LADE) Fu et al. (2024), we show that an adversary can craft malicious inputs that leak prediction hyperparameters like the N-gram size, and the number of prediction candidates generated and verified, which are confidential intellectual property controlling the success rates of a speculative-decoding implementation.
Mitigations. Finally, we discuss potential mitigations against these exploits. We propose transmitting groups of tokens aggregated over several iterations back to the user instead of streaming tokens for each iteration to reduce the observability of fine-grain speculation patterns, and padding packets with a fixed number or random number of tokens to add noise to adversary’s observations to mitigate the fingerprinting attacks. Avoiding usage of private or confidential data for generating predictions also prevents leakage.
@@
    We are the first to observe that variation in the number of tokens generated per autoregressive iteration due to LLM speculative decoding can result in privacy and confidentiality breaches.
@@
    We demonstrate query fingerprinting attacks that can leak out exact matches of private user queries with >90% accuracy and approximate matches with 20% to 40% accuracy.
@@
    We also demonstrate attacks that leak confidential IP controlling the performance of the speculative decoding mechanisms, such as data from data-stores used for prediction or hyper-parameters used for predictions.
@@
    We also propose mitigations including padding packet sizes and aggregating tokens from multiple iterations to limit the potential for these exploits.
2 Background and Motivation
We provide background on speculative decoding in LLMs, and then cover recent side-channel attacks on LLMs.
2.1 Background on Speculative Decoding
Speculative decoding addresses inefficiencies in auto-regressive decoding by allowing multiple token predictions to be verified in parallel, reducing overall latency. It works by generating speculative token guesses, which are then verified by the main model. If correct, multiple tokens are accepted in one step, significantly speeding up the process. If incorrect, the model falls back to standard autoregressive decoding to ensure accuracy.
@@ 2.1 Background on Speculative Decoding
Speculation Using Smaller Models. Many works propose using smaller, much faster draft models to generate sequences of tokens speculatively Miao et al. (2024); Kim et al. (2023), and then verify a single sequence or a tree of such sequences in parallel using the larger target model. The draft model can be a smaller version of the model family or a pruned larger model Yan et al. (2024). In this paper, we use BiLD Kim et al. (2023), as a representative example of this type. Two key policies—fallback and rollback—ensure the quality of tokens is maintained in BiLD: the fallback policy hands control to the larger model if confidence drops, while rollback corrects tokens exceeding a set distance metric, like cross-entropy loss.
Speculation using Self-Drafting. Another speculative decoding paradigm caches previous output tokens and retrieves them for speculation in future iterations, with the target model verifying these speculations Cai et al. (2024); Fu et al. (2024). Look-Ahead Decoding (LADE) Fu et al. (2024) implements this by maintaining a cache where keys are tokens and values are N-grams previously generated after those tokens. LADE populates its cache using jacobi decoding, selecting random tokens from the input and generating token chains. When a key token reappears in the output stream, LADE uses these N-grams for speculation.
@@
Speculation Using Retrieval. Unlike other paradigms relying on model generations, Retrieval-Based Speculative Decoding (REST) He et al. (2024) speculates by retrieving data from a datastore containing contextually relevant text or code. REST’s datastore contains prefixes and continuations from a corpus, matching input suffixes to generate draft outputs. If there are too many drafts, REST selects high-quality subsets and organizes them in a tree, applying a tree attention mask during decoding with the target model.
2.2 Recent Side-Channel Attacks on LLMs
Recent works have demonstrated side-channel vulnerabilities in LLM inference systems leading to critical privacy risks. A recent attack Weiss et al. (2024) demonstrated a token-length side-channel on streaming LLMs, where network-based attackers can leak a large fraction of private response tokens from an LLM chatbot, by observing the size of encrypted packets in the network traffic. By identifying individual tokens based on packet sizes and leveraging inter-sentence context, this attack reconstructs a significant portion of private responses from platforms like ChatGPT and Microsoft Copilot that stream responses token-by-token. Another attack  Debenedetti et al. (2024) demonstrated a tokenizer side-channel, where a malicious user constructs inputs that translate to different number of tokens based on the tokenizer’s vocabulary, which is used to leak out the vocabulary and any private data in the vocabulary.
2.3 Goal: Speculative Decoding Side-Channel
Speculative execution techniques in processors, deployed for improving performance, have resulted in many potent side-channel attacks capable of leaking information, such as Spectre and Meltdown Kocher et al. (2019); Lipp et al. (2018) and their many variants. Consequently, this work seeks to explore whether the deployment of speculative decoding in LLMs similarly for improving inference performance introduces any new side-channels capable of privacy breaches.
@@ 3 Threat Model
Like prior works Weiss et al. (2024), we focus on streaming LLMs, where the user sends requests to an LLM running on the server, and the server sends back the response token-by-token. Here, we study the following attacker models.
  1. 1.
    Network Based Adversary. We assume the attacker is able to observe the network packet size and the time at which they are transmitted, similar to Weiss et al. (2024). The goal of the attacker is to leak private user queries. Although network packets containing LLM response tokens are encrypted, the packet size is still observable. The attacker can also observe the time between each packet to identify which iteration it belongs to. We show this allows an adversary to fingerprint and leak private user queries in Section 4.
  2. 2.
    Malicious User. In this setting, we assume the attacker is a malicious user interacting with the LLM, similar to Debenedetti et al. (2024), with the goal of leaking confidential data used by LLM’s speculative decoding mechanism. Here, the user can craft arbitrary inputs and inspect the output, and also monitor the number of tokens generated per iteration for the output tokens. We show this allows an adversary to leak out any confidential data used for the speculative decoding or its hyperparameters in Section 5.
@@ 4.1 Attack Overview
Attack Scenario. We consider a scenario where a user interacts with an AI chatbot using an LLM. The LLM is programmed to answer a predefined set of user queries through the system prompt. This can occur in scenarios such as a patient in a healthcare setting interacting with an LLM to get medical queries answered, or a user in a customer service setting sharing private information with the chatbot before being redirected to a human. The queries are sent from a client device, over a network, to a server that hosts the LLM equipped with speculative decoding. This leads to the threat of a network-based attacker trying to eavesdrop on the queries and responses, as described in Section 3.
Figure 2: Overview of the Query Fingerprinting Attack. In the offline phase, the network-based attacker profiles the trace of number of tokens generated per iteration influenced by speculative decoding, for all possible inputs, and trains a classifier. In the online phase, the attacker extracts the trace of number of tokens generated per iteration for an unknown input, and uses the trained classifier to leak the input.
Goals and Capabilities. Figure 2 shows the overview of our attack. Our attacker over the network seeks to learn the private query asked by the user. While the query and response packets are encrypted, the attacker can intercept the response packets, measure the time intervals between packets, and count the number of tokens per packet or simply use the packet size to approximate the number of tokens.
Mechanism. Since the input set is limited, the attacker fingerprints inputs based on the pattern of output token counts per iteration influenced by speculative decoding. There are two phases to the attack: offline and online. In the offline phase, the attacker profiles the number of output tokens generated per iteration vs iteration-id (we call this a trace) for each input in the input set, and trains a random forest classifier to predict the input based on the trace. In the online phase, the attacker collects the trace for an unknown input and uses the classifier to identify the input. Next, we provide more details about how these fingerprints are obtained, and then describe how the attack works on various speculative decoding techniques (LADE, REST, BiLD).
@@ 4.2 Fingerprinting Using Speculation Patterns
Speculative decoding introduces variation in the number of tokens generated per iteration. In iterations with correct speculation, multiple tokens are verified in a single iteration and sent back together in a single packet. However, if the speculation is incorrect, a single token is generated per iteration. The attacker uses this variation to create a fingerprint for a particular output sequence.
As shown in  Figure 2, from each packet containing outputs from a single iteration, the attacker measures the number of tokens generated per iteration. Subsequently, by plotting the token numbers vs iteration ID, it gets the trace for an input/output (prompt-response) tuple.
Figure 3 shows the trace of tokens per iteration for two different prompts with LADE across different runs. From this, we observe two key properties that make these traces suitable for usage to identify and leak the prompt and response:
  1. 1.
    Traces are unique to a prompt and response - different prompts result in different traces of tokens per iteration, as shown in Figure 3(a) and Figure 3(b).
  2. 2.
    Traces are reproducible - repeating the same prompt produces similar patterns in the trace of token counts, as shown in the two iterations in Figure 3(a) or the two iterations in Figure 3(b).
(a) What is a narcissistic personality disorder?
(b) What are the symptoms of cancer?
Figure 3: Trace of tokens generated per iteration vs Iteration-ID, for two runs of different prompts (a) and (b) with LADE. The trace has unique and reproducible patterns, enabling its usage as a fingerprint for leaking prompts.
These observations demonstrate that the number of tokens generated per iteration with speculative decoding is a unique identifier for a prompt/response pair. These patterns can leak the prompt/response via side-channels to an adversary.
However, in existing LLM deployments on servers such as OpenAI’s ChatGPT, the logits get decoded into character strings by the detokenization at the server-side before being streamed to the user in encrypted packets. Because the packets are encrypted, the tokens in each packet cannot be directly inferred. Nevertheless, there is a strong correlation between the packet size and the number of tokens per iteration. Figure 4 shows the trace of token count per iteration for a prompt with LADE, and the corresponding packet size in bytes per iteration match considerably. Thus, for our fingerprinting attack, the adversary can use the trace of packet size per iteration as a proxy for the token count per iteration, given that it is easily observable.
Figure 4: Comparing the trace of token count per iteration and trace of packet size, versus iteration-id for a prompt response with LADE. Strong similarity between the traces allows the attacker to infer speculation patterns, i.e., variation in tokens per iteration, via variation in packet size.
Next, we develop attacks to fingerprint and leak user inputs, based on the traces of packet sizes per iteration.
4.3 Attack Experiment Design

### v3

@@ Abstract
Deployed large language models (LLMs) often rely on speculative decoding, a technique that generates and verifies multiple candidate tokens in parallel, to improve throughput and latency. In this work, we reveal a new side-channel whereby input-dependent patterns of correct and incorrect speculations can be inferred by monitoring per-iteration token counts or packet sizes. We demonstrate that an adversary observing these patterns can fingerprint user queries with ¿90% accuracy across four speculative-decoding schemes, REST ($\sim$100%), LADE (up to 92%), BiLD (up to 95%), and EAGLE (up to 77.6%) and leak confidential datastore contents used for prediction at rates exceeding 25 tokens/sec. We evaluate the side-channel attacks in both research prototypes as well as the production-grade vLLM serving framework. To defend against these, we propose and evaluate a suite of mitigations, including packet padding and iteration-wise token aggregation.
@@ 1 Introduction
Large Language Models (LLMs) have transformed natural language processing (NLP), allowing machines to generate and understand human language at an unprecedented scale Vaswani et al. (2017); Devlin et al. (2019); Brown et al. (2020); Zhang et al. (2022); Le Scao et al. (2022); Touvron et al. (2023). LLMs typically generate text using auto-regressive decoding Touvron et al. (2023); OpenAI (2024), where the generation happens serially and each token depends on all the previous ones. Unfortunately, this serial process causes a significant bottleneck in the LLM response latency Miao et al. (2024); Fu et al. (2024) and under-utilizes the available hardware-level parallelism, limiting token generation throughput and latency.
[…]
However, the adoption of speculation techniques is not without risks. Speculative execution in CPUs  Burton (1985); Hennessy & Patterson (2012), which inspired speculative decoding, has led to security vulnerabilities in processors, such as Spectre Kocher et al. (2019) and their variants, which exploit side-channels, i.e. timing variations due to mis-speculations that leak secret data accessed during mis-speculations. This raises the question, does speculative decoding in LLMs also introduce new risks to privacy? In this paper, we provide a study of the privacy risks of speculative decoding in LLMs, including leakage of private user inputs.
@@
Q. How to fine-tune a LLM?   A.  F ine - t uning _a _large _language _model _involves _adjust ing _the _parameters _of _the _model _to _improve _its _performance _on _a _specific _task _or _down stream _task . _Here _are _some _steps _you _can _follow _to _fine - t une _a _large _language
(a) Spec Decoding Example
(b) Tokens Per Iteration
@@
Figure 1(a) shows a response by TinyLlama 1.1B Chat  Zhang et al. (2024a) with a speculative decoding technique, Lookahead Decoding (LADE) Fu et al. (2024), with correct (blue) and mis-speculated (red) tokens illustrated. On correct speculation, up to $n$ tokens are verified per iteration, as shown in Figure 1(b), while on mis-speculation, only one token is generated per iteration. In LLM applications streaming responses to the user, token by token, a network-based adversary, such as a malicious ISP or a compromised router Weiss et al. (2024), can observe network packet sizes corresponding to each iteration to deduce the number of tokens per iteration and learn the input-dependent pattern of correct and incorrect speculations. We target medical chatbots (e.g., Hippocratic AI Shah (2025)), where users interact by asking questions about symptoms or diseases, as this is a privacy-sensitive application where any information leakage (e.g., disease or symptoms of users) breaks privacy laws.
Exploits. We demonstrate privacy breaches based on speculation patterns, as shown in Figure 1(c). First, we show a query fingerprinting attack (e.g., leaking diseases or symptoms in medical chatbots) by a network adversary. By profiling speculation pattern fingerprints for prompt-response pairs offline, an adversary can then use compare the speculation pattern for an unknown query with the fingerprints to recover private user prompts. Across speculative decoding schemes (BiLD, REST, LADE) Fu et al. (2024); He et al. (2024); Kim et al. (2023), our prompt identification attacks reaches accuracies of $\sim$100% for REST, 92% for LADE, and 95% for BiLD with only 30 training samples. On a remote vLLM inference server Kwon et al. (2023), which supports speculation with EAGLE Li et al. (2024a), we similarly observe a high accuracy of 77.6%. We further show that when the actual prompts are unavailable, the adversary can train effective fingerprints using publicly available proxy datasets; e.g., by querying a medical chatbot with the top 50 most common diseases as stand-ins for real prompts. In this case, the attack still achieves 20–40% accuracy in recovering the disease or symptoms from user prompts, far exceeding random guessing (2%). Lastly, we show how adversaries crafting malicious inputs can also leak confidential data such as data-store contents used for predictions with REST He et al. (2024) at a rate of more than 25 tokens/second by observing correctly predicted tokens.
[…]
@@
    We observe that variation in the number of tokens generated per autoregressive iteration due to LLM speculative decoding can result in privacy breaches.
@@
    We demonstrate query fingerprinting attacks that can leak out exact matches of private user queries with >>90% accuracy and approximate matches with 20% to 40% accuracy.
@@
    We also demonstrate attacks that leak confidential IP controlling the performance of the speculative decoding mechanisms, such as data from data-stores used for prediction.
@@
    We propose mitigations including padding packet sizes and aggregating tokens from multiple iterations to limit the potential for these exploits.
2 Related Work
2.1 Prior Speculative Decoding Techniques
Speculative decoding improves the efficiency of auto-regressive decoding in LLMs by verifying multiple token predictions in parallel, reducing latency. These works generate tokens using a smaller model or other heuristics as predictions, which are verified by the target model.
@@ 2.1 Prior Speculative Decoding Techniques
Speculation with Smaller Draft Models. Prior works propose using smaller, faster draft models to generate sequences of tokens speculatively Miao et al. (2024); Kim et al. (2023), and then verify a single sequence or a tree of such sequences in parallel using the larger target model. The draft model can be a smaller version of the model family or a pruned larger model Yan et al. (2025). In this paper, we demonstrate our attacks on BiLD Kim et al. (2023), as a example of this type of speculation.
Speculation via Self-Drafting. Recent works reuse the target model for producing predictions, eliminating the need for a separate draft model. Look-Ahead Decoding (LADE) Fu et al. (2024) caches previous output tokens and uses them to speculate future tokens via a key-value store populated with token sequence N-grams. These are populated using Jacobi decoding, by selecting random tokens from the input and generating successive chain of tokens. When this key token reappears in the output stream, LADE uses the stored N-grams for speculation. Medusa Cai et al. (2024) trains extra heads in the target model, each responsible for a token position in the draft sequence, for speculative tokens. EAGLE and EAGLE-2 Li et al. (2024a; b) perform auto-regressive decoding directly on the feature layers itself to generate speculative tokens, and are integrated in the vLLM inference serving platform. Our attacks target LADE (locally) and EAGLE (on a remote vLLM server) as representative examples, but our techniques are broadly applicable to other techniques as well.
@@
Speculation Using Retrieval. REST He et al. (2024) generates predictions by retrieving data from a datastore containing relevant text or code, organized as a set of prefixes and continuations that can be potentially used as speculative tokens. If there are multiple candidates, REST organizes them in a tree, applying a tree attention mask during decoding with the target model.
Tree-Based Verification. SpecInfer Miao et al. (2024), SpecTr Sun et al. (2023) and REST He et al. (2024) generate multiple drafts of speculation per iteration, organized in a tree structure, to increase the length of successful verification for higher speedup. These drafts are organized with a custom tree attention mask to prevent inter-tree interactions and verified in parallel. In this work, we demonstrate our attacks on REST He et al. (2024), as a representative example of such techniques.
2.2 Side-Channel Attacks on LLMs
Debenedetti et al. (2024) pioneered side-channel attacks on LLMs. Their attack exploits variation in output token counts due to the tokenizer, to leak uncommon strings in the tokenizer vocabulary. Our work instead exploits variation in token count per iteration due to speculative decoding.
Weiss et al. (2024) identified a token-length side channel in streaming LLMs, where network attackers can infer token values by analyzing encrypted packet sizes. This attack reconstructs around 27% of the LLM outputs using the character length of each token. In our work, we similarly focus on a network-attacker observing the encrypted packet sizes, and leverage the variation in packet sizes based on speculative-decoding to identify user inputs. Moreover, we show mitigating the token-length side channel by padding tokens to the maximum token-size still results in packet size variations due to speculative decoding varying the output token counts per packet. We show that this can be used to fingerprint prompts and responses in our attacks with greater than 50% accuracy, even after the token-length side-channel is mitigated, in Appendix D.
Carlini & Nasr (2024) demonstrated timing attacks on LLM inference optimizations in commercial models like GPT-4 and Claude. Their attacks rely on timing variations due to these optimizations, i.e. variations in packet inter-arrival times, which is considerably susceptible to noise due to network congestion wand server loads. In comparison, our attacks exploit packet size variations due to inference optimizations, which are inherently more robust and independent of server or network load variations. We show the resilience of our attacks compared to this prior work in Section 4.7.
Song et al. (2024) and  Zheng et al. (2024) demonstrated timing side-channel attacks exploiting shared key-value (KV) and semantic caches to leak sensitive LLM inputs. Wiretapping LLMs Soleimani et al. (2025) also reveals private attributes in LLM-based medical and financial applications.
@@ 3 Threat Model
Like prior works Weiss et al. (2024), we focus on streaming LLMs, where the LLM server sends back responses to the user token-by-token. Here, we study the following attacker models.
Network Attacker. We assume a malicious ISP or a compromised router that can observe network packet sizes, similar to Weiss et al. (2024). The attacker’s goal is to leak private user queries. Although LLM response packets are encrypted, the packet size is still observable. The attacker can also observe the time between each packet to identify the iteration ID. We show how an adversary can fingerprint and leak private user queries based on this in Section 4.
Malicious User. Here we assume the attacker is a malicious user interacting with the LLM, similar to Debenedetti et al. (2024). The user can craft arbitrary inputs and inspect the output, and the tokens per iteration with the goal of leaking confidential data used by the LLM’s speculative decoding mechanism for generating predictions. We demonstrate this attack in Section 5.
@@ 4.1 Attack Overview
Attack Scenario. Consider a medical AI chatbot, programmed to answer a predefined set of user queries through the system prompt, such as a patient in a healthcare setting asking an LLM medical symptoms related to a disease. The queries are sent from a client device, over a network, to a server that hosts the LLM equipped with speculative decoding. This leads to the threat of a network-based attacker trying to eavesdrop on the queries and responses, as described in Section 3. Figure 2(a) shows the overview of our attack. Our attacker over the network seeks to learn the private query asked by the user. While the query and response packets are encrypted, the attacker can intercept the response packets, and the packet size to approximate the number of tokens per packet.
(a) Overview of the Query Fingerprinting Attack.
(b) Token counts and packet sizes per iteration vs. iteration-id. for a sample response with LADE.
Figure 2: (a) In the offline phase, the network-based attacker profiles variation in number of tokens per iteration influenced by speculative decoding, (based on packet sizes) and trains a classifier. In the online phase, the attacker uses the classifier to leak the input. (b) Packet sizes and tokens per iteration are correlated, allowing variations in packet sizes to be used to approximate token count variations.
Attack Mechanism. Since the input set is limited, the attacker fingerprints inputs based on the pattern of output token counts per iteration influenced by speculative decoding. There are two phases to the attack: offline and online. In the offline phase, the attacker profiles the number of output tokens generated per iteration vs iteration-id (we call this a trace) for each input in the input set, and trains a random forest classifier to predict the input based on the trace. In the online phase, the attacker collects the trace for an unknown input and uses the classifier to identify the input. Next, we provide more details about how these fingerprints are obtained, and then describe how the attack works on various speculative decoding techniques (LADE, REST, BiLD, and EAGLE).
@@ 4.2 Fingerprinting Using Speculation Patterns
Speculative decoding causes variation in the number of tokens generated per iteration, as correct speculation results in multiple tokens verified and sent together, while incorrect speculation generates one token per iteration. This variation can be used to fingerprint prompt/response pairs. We observe that the variations in tokens per iteration versus iteration-ID are both unique to each prompt-response combination, and are also reproducible to a great extent, as shown in Figure 4 in Appendix B.
To quantitatively validate this, we generated 2 traces each from 50 prompts taken from MedAlpaca dataset Han et al. (2023) (Human-LLM interactions asking questions on different diseases; see Appendix A) input to LADE, with temperature of 0.3, and computed the pairwise cosine similarity between them. Traces of the same prompt (or disease) have cosine similarity of 0.9–1, whereas traces from different prompts have cosine similarity between 0.4 - 0.8. This shows that the trace of tokens per iteration with speculative decoding can uniquely fingerprint and identify a prompt and response.
As packets are encrypted, the attacker cannot see the token counts directly. However, as shown in Figure 2(b), packet size strongly correlates with the number of tokens per iteration (Pearson Correlation Coefficient of 0.747 across all packets above). Thus, encrypted packet sizes serves as an effective proxy for the tokens per iteration in our fingerprinting attack. Note that exact token counts are not necessary; merely correlation of packet size with token counts suffices. Furthermore, any mitigation for prior token-size side-channels Weiss et al. (2024) by hiding individual token sizes is insufficient, as speculation itself introduces packet size variations due to varying token counts (see Appendix D).
4.3 Attack Design